OpenStudy (moses238):

Charlies Car Rentals charges a flat fee of $20 plus $24/day to rent a car. Jerrys Car Rentals charges $28/day with no flat fee. The system that models this situation is given, where c is the cost of renting a car and d is the number of days a car is rented. c = 20 + 24d c = 28d The solution to the system is (5, 140). Which interpretation correctly describes the solution to the system of equations? A. The cost to rent a car is the same, $140, for both car rental companies if you rent a car for 5 days. B. The cost to rent a car is the same, $5, for both car rental companies if you rent a car for 140 days. C. Company A will charge more money on the fifth day by charging $140. D. Company A will charge less money on the fifth day by charging $140.

OpenStudy (texaschic101):

A. subbing in 5 for d (number of days)...we get 20 + 24(5) = 20 + 120 = 140 28(5) = 140